OpenStudy (anonymous):

Which value of x makes the equation true? x – 5.1 = –7.6 x =

OpenStudy (chrisplusian):

-2.5

OpenStudy (anonymous):

i need your work

OpenStudy (chrisplusian):

Ok hang on one second I am slow with the equation editor

OpenStudy (chrisplusian):

\[x-5.1=-7.6\] Now add 5.1 to both sides ...

OpenStudy (chrisplusian):

\[x-5.1+5.1=-7.6+5.1\]

OpenStudy (chrisplusian):

\[x+0=-2.5\]

OpenStudy (chrisplusian):

\[x=-2.5\]

OpenStudy (chrisplusian):

Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (chrisplusian):

Cool